Assume a demand equation Qd # 9-01p-pc + 001ps +00001Y and a supply equation: Qs = 0.1p-002pī + 001N + 0.01T-01w where Q quantity in thousands of units p=price of the good Pc npnce of a complement - S3P Ps price of a substtute $200 Y tt consumer income = S80.000 pnice of an input N number of tirms700 $450 T - index of technology we wage rate 300 y wage rate - s40 r the price is $105, there will be an orthousand units
Answer #1

Qd = 9 - 0.1p - pc + 0.01ps + 0.0001Y = 9 - 0.1(105) - 3 + 0.01(200) + 0.0001(80,000) = 9 - 10.5 -3 +2+8 = 5.5

Qs = 0.1p - 0.02pi + 0.01N + 0.01T - 0.1w = 0.1(105) - 0.02(450) + 0.01(700) +0.01 (300) - 0.1(40) = 10.5 -9 +7 +3 -4 = 7.5

Qs - Qd = 2

Therefore, at price $105, there will be an excess supply of 2 thousand units